OpenStudy (anonymous):

Question: 86 of 95: Evaluate the expression for x = 4 and y = 5. 6(x2 - y) - y2 Select one of the options below as your answer:A. 8B. 41C. 47D. 56

OpenStudy (raden):

subtitute the values of x=4 and y=5, it means 6(4^2 - 5) - 5^2 = ...

OpenStudy (anonymous):

which letter?

OpenStudy (raden):

u first

OpenStudy (anonymous):

B?

OpenStudy (anonymous):

41 B?

OpenStudy (raden):

6(4^2 - 5) - 5^2 = 6(16-5) - 25 = 6(11) - 25 = 66-25 = 41 u are right :)

OpenStudy (anonymous):

:D